# Novel View Synthesis using DDIM Inversion

### v1

@@ Introduction
Fusion Strategy
To synthesize semantically rich target view latents from the predicted DDIM-inverted mean latent $\tilde{\mathbf{z}}^{\text{inv}}_{\text{tar},\mu}$, we introduce two fusion strategies that combine this mean latent with a noise component derived from the input view latent. These strategies re‑inject the learned noise variance, that is, the high‑frequency details, into the coarse latent. We utilize the fact that the noise/variance term Equation (2) in the DDIM-inverted latent of the input view contains scene-level attributes and characteristics (Staniszewski, Kuciński, and Deja 2025), which can be used to synthesize the scene from a novel view when fused with TUNet’s prediction.
Strategy A: Variance Fusion via $\sigma$-Component
@@
The fused latent $\mathbf{z}_{\text{noisy}}$ is then passed into the Stable Diffusion U-Net to compute the noise prediction:
$\boldsymbol{\epsilon}_{\theta}=\text{U-Net}(\mathbf{z}_{\text{noisy}},t).$ (5)
The initial latent for DDIM sampling is obtained as:
$\tilde{\mathbf{z}}_{\text{tar}}^{\text{inv}}=\tilde{\mathbf{z}}_{\text{tar},\mu}^{\text{inv}}+\sqrt{1+\alpha_{t+1}}\boldsymbol{\epsilon}_{\theta}.$ (6)
Strategy B: Direct Noise Addition from Reference Inversion
@@
$\tilde{\mathbf{z}}_{\text{tar}}^{\text{inv}}=\tilde{\mathbf{z}}_{\text{tar},\mu}^{\text{inv}}+\sqrt{1+\alpha_{t+1}}\mathbf{z}_{\text{ref}}^{\text{inv}}.$ (7)
We generate samples using $\tilde{\mathbf{z}}_{\text{tar}}^{\text{inv}}$ from both Equation (6) and Equation (7), and pick the best generated result where we pick the result which has lower LPIPS score.
@@
Experiments
Dataset: MvImgNet (Yu et al. 2023b) consists of 6.5 million frames of real-world scenes across 238 categories. We use two subsets of MVImgNet (i) three scene categories: sofas, chairs, and tables. A 90-5-5 (Anciukevičius et al. 2024) split is used for training, validation, and testing, respectively, determined by lexicographic ordering of the scene identifiers. (ii) We use 8.5 lakh frames across 167 classes and for each class, we keep 1 scene out of 99 in the test set to evaluate our results and compare with other methods.
@@
Implementation Details : TUnet has approximately 148M parameters. The dimensions of both class and camera embeddings are 64, and the cross-attention dimension is 768 with an attention head dimension of 64. We use the latent diffusion backbone (Rombach et al. 2022). For training, we randomly pair frames 1-10 of each scene with frames 15-25. We effectively use 20 frames per scene for training. We adopt the same frame pairing strategy for our evaluation. We train two models on our subset (i) 3 classes and (ii) 167 classes. Our 167 class model is trained for 450 epochs on a single 40GB A400 for 17 GPU days with a batch size of 32 and a learning rate of 1e-5, and we decay the learning rate using a cycle scheduler. During inference, we generate final results with 30 DDIM sampling steps with the initial latent being Equation (6) or Equation (7), which represents the noisy latent at $t$ = 600.
We compare our 3-class model with GIBR (Anciukevičius et al. 2024) and the 167-class model with NViST (Jang and Agapito 2024). For GIBR, we have the same train/test split and directly report the results from their paper. For NViST, we use their pre-trained model to test exact input/target frame pairs. All of the testing frames are from unseen scenes within the classes used in training. To compare with GIBR, we resize our $512\times 512$ results to $256\times 256$. For direct comparison with NViST, we resize our results to $90\times 90$. We report LPIPS and FID scores with the resized results for comparisons. We report PSNR and SSIM with the original $512\times 512$ resolution results.
Quantitative Comparison
@@
Qualitative Comparison
We compare our results with NViST (Jang and Agapito 2024) in Figure 5. It is evident that our method synthesizes the target with high fidelity and is able to generate results for near as well as far target views, where NViST fails.
@@
Unseen classes: We show qualitative results on 6 unseen classes in Figure 6. We cover outdoor and indoor scenes, as well as include large and small object classes in the test set. Even on unseen classes, we are able to generate high-resolution reconstruction for a diverse set of scenes.
@@
Out of domain data: To evaluate zero‑shot generalization beyond MVImgNet, we assembled an out‑of‑domain test set by downloading freely‑licensed photographs from Unsplash (Unsplash Contributors 2025) featuring natural scenes. We show the results and compare with Zero123++ (Shi et al. 2023) in Figure 7. While both methods successfully generate plausible novel viewpoints, our approach produces more faithful surface textures and preserves natural scene characteristics.
@@
Ablation
The ablation results are presented in Table 3. We evaluate the following settings. In the first setting, Concat, we concatenate class and camera embeddings with the input, but do not inject them into every ResNet block. Here, we see that there is a significant performance drop. Second, w/o cross-attn, where we remove all cross-attention layers. The results degrade for all three metrics.
@@
Table 3: Ablation study using 3-class.
Conclusion
@@
Novel View Synthesis using DDIM Inversion
Supplementary Materials
Novel View Synthesis and Editing
As we work with DDIM inverted noise, we have the flexibility to extend our pipeline to editing via textual prompts during the final sampling. For our novel view reconstruction, we use BLIP (Li et al. 2022) generated captions for the input image $\mathbf{x}_{ref}$. We edit these prompts to add desired scenarios in our reconstructed novel view. Figure 8 shows results of our editing mechanism.
Figure 8: We use textual prompts to edit the novel view reconstructions as desired. In the first row, we place a bag on the bench. In the second row, a dog is introduced in the scene. In the third row, a pair of shoes are placed by the bench.
Synthesis of Multiple Views from Single Image
@@
Figure 9: Generating multiple frames with single input image.
Qualitative Ablation Results
@@
Figure 10: Ablation
Additional Results
@@
We show additional results using our 3-class and 167-class models in Figures 11 and 12, respectively.

### v2

@@ 3.3.4 Cross-Attention Module
3.4 Fusion Strategy
To synthesize semantically rich target view latents from the predicted DDIM-inverted mean latent $\tilde{\mathbf{z}}^{\text{inv}}_{\text{tar},\mu}$, we introduce two fusion strategies that combine this mean latent with a noise component derived from the input view latent. These strategies re‑inject the learned noise variance, that is, the high‑frequency details, into the coarse latent. We utilize the fact that the noise/variance term Equation (2) in the DDIM-inverted latent of the input view contains scene-level attributes and characteristics [41], which can be used to synthesize the scene from a novel view when fused with TUNet’s prediction.
3.4.1 Strategy A: Variance Fusion via $\sigma$-Component
@@
The fused latent $\mathbf{z}_{\text{noisy}}$ is then passed into the Stable Diffusion U-Net to compute the noise prediction, $\boldsymbol{\epsilon}_{\theta}=\text{U-Net}(\mathbf{z}_{\text{noisy}},t).$ The initial latent for DDIM sampling is obtained as:
$\tilde{\mathbf{z}}_{\text{tar}}^{\text{inv}}=\tilde{\mathbf{z}}_{\text{tar},\mu}^{\text{inv}}+\sqrt{1+\overline{\alpha}_{t+1}}\boldsymbol{\epsilon}_{\theta}.$ (5)
3.4.2 Strategy B: Direct Noise Addition from Reference Inversion
@@
$\tilde{\mathbf{z}}_{\text{tar}}^{\text{inv}}=\tilde{\mathbf{z}}_{\text{tar},\mu}^{\text{inv}}+\sqrt{1+\overline{\alpha}_{t+1}}\mathbf{z}_{\text{ref}}^{\text{inv}}.$ (6)
We generate samples using $\tilde{\mathbf{z}}_{\text{tar}}^{\text{inv}}$ from both Equation (5) and Equation (6).
@@ 3.5 Training Objective
4 Experiments
Dataset: We perform experiments using MvImgNet [61] and RealEstate10K [66]. MvImgNet consists of 6.5 million frames of real-world scenes across 238 categories. We use two subsets of MVImgNet (i) three scene categories: sofas, chairs, and tables. A 90-5-5 [1] split is used for training, validation, and testing, respectively, determined by lexicographic ordering of the scene identifiers. (ii) We use 8.5 lakh frames across 167 classes and for each class, we keep 1 scene out of 99 in the test set to evaluate our results and compare with other methods. In case of RealEstate10K, we train using 1 million pairs. We report additional results of RealEstate10K in the supplementary and demonstrate that our method achieves superior results.
@@ 4 Experiments
Implementation Details : TUnet has approximately 148M parameters. The dimensions of both class and camera embeddings are 64, and the cross-attention dimension is 768 with an attention head dimension of 64. We use the latent diffusion backbone [35]. For training, we randomly pair frames 1-10 of each scene with frames 15-25. We effectively use 20 frames per scene for training. We adopt the same frame pairing strategy for our evaluation. We train two models on our subset (i) 3 classes and (ii) 167 classes. Our 167 class model is trained for 450 epochs on a single 49 GB RTX A6000 for 17 GPU days with a batch size of 32 and a learning rate of 1e-5, and we decay the learning rate using a cycle scheduler. During inference, we generate final results with 30 DDIM sampling steps with the initial latent being Equation (5) or Equation (6), which represents the noisy latent at $t$ = 600.
We compare our 3-class model with GIBR [1] and the 167-class model with NViST [22]. For GIBR, we have the same train/test split and directly report the results from their paper. For NViST, we use their pre-trained model to test exact input/target frame pairs. All of the testing frames are from unseen scenes within the classes used in training. To compare with GIBR, we resize our $512\times 512$ results to $256\times 256$. For direct comparison with NViST, we resize our results to $90\times 90$. We report LPIPS and FID scores with the resized results for comparisons. We follow the evaluation protocol as given in [1, 8].
4.1 Quantitative Comparison
@@
4.2 Qualitative Comparison
We compare our results with NViST [22] in Figure 5. It is evident that our method synthesizes the target with high fidelity and is able to generate results for near as well as far target views, where NViST fails.
@@ 4.2 Qualitative Comparison
Unseen classes: We show qualitative results on 6 unseen classes in Figure 6. We cover outdoor and indoor scenes, as well as include large and small object classes in the test set. Even on unseen classes, we are able to generate high-resolution reconstruction for a diverse set of scenes. For evaluation on unseen classes, the unseen class is treated as an additional label. We obtain its semantic embedding and provide it to the model at test time.
@@
Out of domain data: To evaluate zero‑shot generalization beyond MVImgNet, we assembled an out‑of‑domain test set by downloading freely‑licensed photographs from Unsplash [47] featuring natural scenes. Since web images lack ground-truth camera parameters, we identify the most visually similar scene in MvImgNet in terms of viewpoint. The camera parameters of this nearest neighbor are then used as a proxy for the web image. For target views, we analogously select the corresponding frame from the same scene in which the closest viewing-angle image resides, and adopt its parameters. We show the results and compare with Zero123++ [39] in Figure 7. While both methods successfully generate plausible novel viewpoints, our approach produces more faithful surface textures and preserves natural scene characteristics.
@@
4.3 Ablation Study
Architecture Design: The model design ablation results are presented in Table 3. We evaluate the following settings. In the first setting, Concat, we concatenate class and camera embeddings with the input, but do not inject them into every ResNet block. Here, we see that there is a significant performance drop. Second, w/o cross-attn, where we remove all cross-attention layers. The results degrade for all three metrics.
@@ 4.3 Ablation Study
Table 3: Ablation study using 3-class model.
Fusion Strategy Comparison: We compare perceptual and image quality assessment metrics for the two fusion strategies in Table 4. Variance Fusion achieves better results in all metrics.
Fusion Method LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$
Variance (Stgy A) 0.495 15.41 0.521 69.86
Direct (Stgy B) 0.521 14.76 0.457 102.72
Table 4: Comparison of perceptual and image-quality assessment metrics between the two fusion strategies, using 3-class model.
Different Stable Diffusion Pipelines: We evaluate our pipeline using three diffusion backbones. Stable Diffusion v1.5 [35] is the pipeline we use by default in all of our experiments. Furthermore, we compare our default pipeline with v2.1111https://huggingface.co/stabilityai/stable-diffusion-2-1, and Dreamlike Photoreal 2.0222https://huggingface.co/dreamlike-art/dreamlike-photoreal-2.0. As shown in Table 5, performance remains consistent across models, indicating robustness of our framework. v2.1 achieves slightly lower FID, reflecting improved generative realism, while v1.5 attains marginally higher PSNR/SSIM. Dreamlike Photoreal 2.0 shows increased artifacts, likely due to its photoreal art domain fine-tuning, which makes it perform worse for natural images.
Pipeline LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$
Stable Diffusion v1.5 0.491 15.41 0.522 93.57
Stable Diffusion 2.1 0.491 15.27 0.522 88.95
Dreamlike Photoreal 2.0 0.507 15.15 0.503 101.84
Table 5: Comparison across different SD pipelines.
Different Diffusion Timesteps: We compare the performance at three different timesteps $t=400,600,800$. As shown in Table 6, decreasing the diffusion timestep from $t=600$ to $t=400$ leads to a degraded performance as reflected in the scores. In our experiments, we observe that at $t=400$, the loss is very high compared to the case of $t=600$. This indicates that training is harder for a single-step translation with TUnet for $t=400$. At $t=400$, the noise level and the inversion trajectory are insufficient for meaningful variance-based fusion. The latent still retains the dominant low-frequency structure, reducing the fusion strategy’s ability to recover and refine high-frequency content. The performance is worst at $t=800$ as the inversion at this timestep loses most of its signal to perform any effective translation. In contrast, $t=600$ allows a weak yet sufficient signal for effective view translation aided by noise fusion to recover better quality results upon sampling.
Timestep ($t$) LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$
400 0.510 15.37 0.528 150.28
600 0.491 15.41 0.522 93.57
800 0.550 15.01 0.502 197.13
Table 6: Comparison at different diffusion timesteps $t$.
4.4 RealEstate10K
We compare our results with GenWarp [38] and VIVID [10]. We use 1K images for testing. In Table 1, we can see that our method performs better in terms of LPIPS, PSNR, and SSIM, except for long range LPIPS compared to VIVID.
Mid-range (30-60 frames) Long-range (60-120 frames)
Method LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
VIVID 0.523 13.83 0.439 0.594 12.69 0.410
Ours 0.503 15.04 0.479 0.609 13.44 0.448
Table 7: Results on 1K pairs of RealEstate10K. Images are uniformly sampled at random from different scenes.
5 Conclusion
@@
1 Additional Experimental Results
We provide additional results on RealEstate10K and MVImgNet, including qualitative ablations and high-resolution evaluations. We further study task-level generalization by applying the same DDIM-latent translation and fusion principle to low-light image enhancement, treating it as an image-to-image translation problem.
1.1 RealEstate10K
In order to train the network, we sample 1 million source-target pairs. We resize the images while preserving the aspect ratio, and then center-crop a region of $\times256$. Extrinsics parameters remain the same for all image resolutions, following VIVID. To obtain intrinsics, we selected the focal length and principal point based on a resolution of 256 $\times$ 256. As RealEstate10K do not have class labels, we do not use class embeddings. We follow the evaluation pipeline of VIVID. We compute the metrics for $\times256$ resolution. To generate the images from GeoGPT, Photometric-NVS, and VIVID, we follow the respective sampling and pre-processing strategy as mentioned in the paper.
We evaluate on 1K mid-range (30–60 frames) and 1K long-range (60–120 frames) RealEstate10K test pairs. All metrics are reported as mean $\pm$ standard deviation over three independently sampled subsets of the official test split. We report the results in Table 1. While our LPIPS is higher than prior work, we improve PSNR and SSIM over baselines Photometric-NVS and VIVID, especially on long-range pairs where viewpoint extrapolation is most challenging. Our method shows (i) best long-range SSIM, (ii) strong PSNR gains over Photometric-NVS and VIVID, and (iii) low variance across sampled subsets.
Mid-range (30-60 frames) Long-range (60-120 frames)
Method LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
GeoGPT 0.318 $\pm$ 0.004 15.863 $\pm$ 0.097 0.474 $\pm$ 0.008 0.409 $\pm$ 0.005 14.025 $\pm$ 0.223 0.416 $\pm$ 0.016
Photometric-NVS 0.387 $\pm$ 0.004 14.271 $\pm$ 0.023 0.410 $\pm$ 0.002 0.508 $\pm$ 0.001 12.328 $\pm$ 0.086 0.342 $\pm$ 0.009
VIVID 0.442 $\pm$ 0.070 13.480 $\pm$ 0.309 0.408 $\pm$ 0.027 0.547 $\pm$ 0.040 11.852 $\pm$ 0.733 0.340 $\pm$ 0.060
Ours 0.510 $\pm$ 0.008 14.902 $\pm$ 0.026 0.471 $\pm$ 0.003 0.609 $\pm$ 0.002 13.243 $\pm$ 0.052 0.430 $\pm$ 0.005
Table 1: Results on RealEstate10K.
[…]
Method #Parameters (M) Inference time ($\times$ Ours)
[…]
Table 2: Model size and inference speed comparison on RealEstate10K (relative to Ours). Our method takes 2 seconds.
In Figure 1, we show the output generated by different methods. Compared to prior work, our method produces slightly softer textures but preserves global scene structure and viewpoint alignment. This is consistent with our quantitative metrics: despite higher LPIPS, we obtain PSNR/SSIM improvements on challenging long-range pairs, with a compact model that remains efficient at inference.
Input Target VIVID PhotoNVS GeoGPT Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 1: Qualitative comparison on RealEstate10K.
1.2 Synthesis of Multiple Views from Single Image
@@
1.3 Qualitative Ablation Results
@@
1.4 Additional Results
@@
We show additional results using our 3-class and 167-class models in Figures 4 and 5, respectively. In Figure 4, our model produces novel views that remain geometrically consistent with the targets while preserving global scene layout. Figure 5 further shows improved texture fidelity and sharper high-frequency details.
1.5 Failure Cases
In Figure 6, we show the failure cases. In the first column, we can see that structure is distorted. In the second and third columns, shape is not preserved. Similarly, for other columns, we see that texture, count, or shape is not preserved.
2 Extension to Low Light Image Enhancement Task
In this section, we show that our method can be extended to image-to-image translation task such as LLIE. As LLIE requires high fidelity with respect to input-target pairs, we make use of depth maps. The depth maps for input images are obtained using Intel DPT-Large model [^3] https://huggingface.co/Intel/dpt-large. As the camera parameters are not available, we do not employ them.
We evaluate the LLIE application of our proposed method on multiple datasets, including LOLv1 [53], LOLv2 [58], and SICE [4]. LOLv1 comprises 485 paired low-light and normal-light training images and 15 testing pairs. LOLv2 is divided into LOLv2-Real and LOLv2-Synthetic subsets, each containing 689 and 900 training pairs and 100 testing pairs, respectively. SICE [4] contains 589 low-light and overexposed images. We use 80% of scenes for training and 20% for testing. The training image resolution is $\times256$.
Experiment Settings The diffusion process employs a linear noise schedule with $\beta_{\text{start}}=0.0001$ and $\beta_{\text{end}}=0.02$ over $T=500$ timesteps. The model is optimized using AdamW optimizer with learning rate $1\times 10^{-4}$, batch size 16, and trained for 1000 epochs using cosine annealing learning rate scheduling. The loss function consists of MSE and LPIPS [64] with weighting factor $\lambda=0.1$ for LPIPS loss. Depth conditioning is achieved by concatenating 4-channel low-light latents with single-channel depth maps estimated from enhanced or ground truth images. At the test time, we use CIDNet [56] to obtain enhanced images which are used to further obtain depth maps. As the number of training images is less, we train using two different settings. In the first setting, we combine LOLv1 and LOLv2 dataset. As LOLv2 has LOLv2-Synthetic dataset, in order to train only on real datasets, we use a second setting wherein we combine LOLv1, LOLv2-Real, and SICE datasets.
Evaluation Metrics For quantitative evaluation, we adopt Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) [52] as distortion metrics. To assess perceptual quality, we report Learned Perceptual Image Patch Similarity (LPIPS) [64] with AlexNet [25] backbone. Model evaluation is conducted using DDIM sampling with 50 denoising steps.
Main Results Table 6 presents the quantitative comparison of our method against state-of-the-art low-light enhancement techniques. In case of LOLv2, our method outperforms all the other methods by a huge margin. As LOLv2 is larger compared to LOLv1 and a real dataset, the improvement on this dataset is extremely significant. In case of LOLv1, our method shows the best LPIPS score.
To ensure a fair comparison, we also train CIDNet [56] and RetinexFormer [6] using the combined dataset. We present these results in Table 6. In case of CIDNet, we see that the performance significantly drops in terms of PSNR and LPIPS, though it shows marginal improvement in SSIM. The drop in the performance is more pronounced in LOLv2-Synthetic. We observe a similar phenomenon in case of RetinexFormer. The performance, however, improves in case of LOLv2-Real for both CIDNet and RetinexFormer.
We further experiment with only real datasets and report results in Table 6. We combine LOLV1, LOLV2-Real and SICE. We see that for both LOLv2-Real and SICE, our method shows best performance in terms of PSNR. In LOLv1, CIDNet performs best.
We show a visual comparison of our model output with RetinexFormer and CIDNet in Figure 7. Even though our method is not explicitly designed for LLIE task, we can see that our model generalizes very well to this task.
Figure 2: Generating multiple frames with single input image from MVImgNet.
Figure 3: Qualitative ablation results.
@@ 2 Extension to Low Light Image Enhancement Task
Input Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Novel Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Target Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: MVImgNet failure cases. Each column represents one scene, while rows correspond to input view, predicted novel view, and target view.
Methods Color Model Complexity LOLv1 LOLv2-Real LOLv2-Synthetic
[…]
SNR-Aware [55] SNR+RGB 4.01 26.35 26.716 0.851 0.152 21.480 0.849 0.163 24.140 0.928 0.056
Bread [16] YCbCr 2.02 19.85 25.299 0.847 0.155 20.830 0.847 0.174 17.630 0.919 0.091
PairLIE [13] Retinex 0.33 20.81 23.526 0.755 0.248 19.885 0.778 0.317 19.074 0.794 0.230
LLFormer [50] RGB 24.55 22.52 25.758 0.823 0.167 20.056 0.792 0.211 24.038 0.909 0.066
RetinexFormer [6] Retinex 1.53 15.85 27.140 0.850 0.129 22.794 0.840 0.171 25.670 0.930 0.059
GSAD [20] RGB 17.36 442.02 27.605 0.876 0.092 20.153 0.846 0.113 24.472 0.929 0.051
CIDNet [56] HVI 1.88 7.57 28.201 0.889 0.079 24.111 0.871 0.108 25.705 0.942 0.045
Ours RGB 182.04 8.05 27.358 0.848 0.075 28.097 0.810 0.059 19.663 0.671 0.140
Table 4: LOLv1 and LOLv2 results. Following CIDNet [56], we use GT mean method during testing for LOLv1. Best performance is in bold.
Methods Color Model LOLv1 LOLv2-Real LOLv2-Synthetic
[…]
CIDNet HVI 23.103 0.902 0.106 29.455 0.927 0.071 17.462 0.851 0.201
RetinexFormer Retinex 22.844 0.827 0.146 28.400 0.877 0.116 16.032 0.749 0.254
Ours RGB 27.358 0.848 0.075 28.097 0.810 0.059 19.663 0.671 0.140
Table 5: Results for combined dataset. represents the training on the combined dataset.
Methods Color Model LOLv1 LOLv2-Real SICE
PSNR SSIM PSNR SSIM PSNR SSIM
RUAS [34] Retinex 18.654 0.518 15.326 0.488 8.656 0.494
LLFlow [51] RGB 24.998 0.871 17.433 0.831 12.737 0.617
CIDNet [56] HVI 28.201 0.889 24.111 0.871 13.435 0.642
CIDNet∗∗ HVI 20.560 0.808 22.109 0.841 13.227 0.378
Ours RGB 26.859 0.845 28.182 0.807 15.554 0.382
Ours HVI 26.753 0.843 28.072 0.805 15.565 0.390
Table 6: Results on LOLv1, LOLv2-Real, and SICE datasets. ∗∗ represents the training on the combined dataset.
Input RetinexFormer RetinexFormer RetinexFormer∗∗ CIDNet CIDNet
CIDNet∗∗ Ours Ours∗∗ Ours-HVI∗∗ Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input RetinexFormer RetinexFormer RetinexFormer∗∗ CIDNet CIDNet
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
CIDNet∗∗ Ours Ours∗∗ Ours-HVI∗∗ Ground Truth
Figure 7: Visual comparison of RetinexFormer, CIDNet, and our method across training configurations. Ours, indicates training using LOLv1+LOLv2; ∗∗ indicates training using LOLv1+LOLv2+SICE; Ours-HVI uses HVI color model.